### v1

@@ 4 Conclusions
We analytically show that the limited expressivity of QGAN models stems from a fundamental constraint: a lower bound on the fidelity between the generator’s pure-state output and the target data distribution. This bound limits the generator’s ability to match complex data, thereby conflicting with the optimal training of the quantum discriminator.
Consequently, quantum generative models that rely solely on pure-state outputs—without post-selection or sampling—face intrinsic limitations when modelling high-rank (i.e., mixed or complex) data distributions. In such cases, the generator cannot truly generalize; instead, it approximates a fixed representative of the data, effectively reducing the generative task to deterministic reproduction.
@@ Acknowledgements
Funding
This work was funded by the Quanten-Initiative Rheinland-Pfalz (QUIP) and Research Initiative Quantum Computing for AI (QC-AI).
@@ Appendix A Loss Plots
Appendix B Limitations of a QGAN Generator
@@
As illustrated in the schematic of Fig. 5(a), the layers of the classical NN are not fully connected. On the contrary, each neuron is connected only to its corresponding neuron in the next layer. Additionally, the weights are fixed and set to 1, ensuring that the input values remain unchanged. The only trainable parameters are the biases, which are added to the inputs before being passed through a non-linear activation function applied to the hidden layer neurons. This deliberately constrained architecture is intended to isolate and examine the effect of non-linearity. As for the QuGAN generator—one layer of which is illustrated in Fig. 5(a)—we set the number of layers such that the total number of trainable parameters is approximately equal to that of the classical neural network (i.e. the number of nodes). However, we acknowledge that this does not constitute a strictly fair comparison, as the representational capacities and training dynamics of quantum and classical models differ significantly. In Fig. 5(c)(i) we present the target image. Adjacent to it, Fig. 5(c)(ii) shows of the classical neural network, and Fig. 5(c)(iii) displays the output of the QVC. We observe that even this highly constrained version of a classical neural network is able to closely approximate the target image, whereas the QVC clearly struggles to do so. A common approach to inducing non-linear transformations in quantum circuits involves introducing ancillary qubits, performing partial measurements, and tracing out the ancillary subsystem [26, 50]. The non-linear effect can be amplified by increasing the number of ancillary qubits. As shown in Fig. 5(c)(iv), the generated output begins to approximate the target image more closely as the number of ancillary qubits approaches that of the generator’s qubits (i.e., 12 qubits in total, with 6 ancillary). We also have to note that in this case the resulting generated state is not pure, therefore the limitations presented in Sec. 7 do not apply. In this case, once the model is trained, providing noise vectors as input leads to the generation of output states that exhibit a degree of variability (see Fig. 5(c)(iv)). However, our attempts to train generative variational circuits with the aforementioned characteristics on multiple classes of the MNIST dataset were unsuccessful.

### v2

@@ 4 Conclusions
Beyond our numerical analysis of existing QGAN models, a central contribution of this work is the explicit analytical connection we establish between optimal discriminator performance and the eigenvalue structure of the data density matrix when the generator outputs a single pure quantum state. Namely, we analytically show that the limited expressivity of QGAN models stems from a fundamental constraint: a lower bound on the fidelity between the generator’s pure-state output and the target data distribution. This bound limits the generator’s ability to match complex data, thereby conflicting with the optimal training of the quantum discriminator.
This result complements existing studies on the expressivity of quantum generative models (e.g. QGANs [24, 26, 36], PQCs [60, 61], and QCBMs [62]) by highlighting that approaches relying solely on pure-state outputs—without post-selection or sampling—face intrinsic limitations when modelling high-rank (i.e., mixed or complex) data distributions. In such cases, the generator cannot truly generalize; instead, it approximates a fixed representative of the data, effectively reducing the generative task to deterministic reproduction.
While our numerical simulations focus on classical datasets encoded into mixed quantum states, the analytical results presented in this paper are not specific to classical data generation and apply equally to the generation of quantum data. Our analysis is formulated entirely at the level of quantum states and quantum distinguishability measures, independent of whether the target state arises from classical data encoding or represents an intrinsically quantum dataset.
@@ Appendix A Loss Plots
Appendix B Training details
All models are trained using the Adam optimizer with a learning rate of 0.001 and a batch size of 32. Training is performed for 10 epochs, which we keep fixed since the generator loss and fidelity empirically converge within this range, with no noticeable improvement beyond that. For the QGAN setup, both the generator and discriminator are parametrized quantum circuits operating on $k$ qubits, where $k$ corresponds to the number of retained PCA components. Per iteration, the discriminator is trained using a real-data minibatch and 16 fake samples generated by the generator, while the generator is subsequently updated using 8 discriminator evaluations. After each epoch, a sample is generated by measuring the generator qubits, forming a $k$-dimensional vector that is mapped back to an image via inverse normalization and inverse PCA.
Appendix C Limitations of a QGAN Generator
@@
As illustrated in the schematic of Fig. 6(a), the layers of the classical NN are not fully connected. On the contrary, each neuron is connected only to its corresponding neuron in the next layer. Additionally, the weights are fixed and set to 1, ensuring that the input values remain unchanged. The only trainable parameters are the biases, which are added to the inputs before being passed through a non-linear activation function applied to the hidden layer neurons. This deliberately constrained architecture is intended to isolate and examine the effect of non-linearity. As for the QuGAN generator—one layer of which is illustrated in Fig. 6(a)—we set the number of layers such that the total number of trainable parameters is approximately equal to that of the classical neural network (i.e. the number of nodes). However, we acknowledge that this does not constitute a strictly fair comparison, as the representational capacities and training dynamics of quantum and classical models differ significantly. In Fig. 6(c)(i) we present the target image. Adjacent to it, Fig. 6(c)(ii) shows of the classical neural network, and Fig. 6(c)(iii) displays the output of the QVC. We observe that even this highly constrained version of a classical neural network is able to closely approximate the target image, whereas the QVC clearly struggles to do so. A common approach to inducing non-linear transformations in quantum circuits involves introducing ancillary qubits, performing partial measurements, and tracing out the ancillary subsystem [31, 59]. The non-linear effect can be amplified by increasing the number of ancillary qubits. As shown in Fig. 6(c)(iv), the generated output begins to approximate the target image more closely as the number of ancillary qubits approaches that of the generator’s qubits (i.e., 12 qubits in total, with 6 ancillary). We also have to note that in this case the resulting generated state is not pure, therefore the limitations presented in Section “Fidelity Bounds in QGANs” do not apply. In this case, once the model is trained, providing noise vectors as input leads to the generation of output states that exhibit a degree of variability (see Fig. 6(c)(iv)). However, our attempts to train generative variational circuits with the aforementioned characteristics on multiple classes of the MNIST dataset were unsuccessful.